# Small-Signal Model for Inhomogeneous Helix Traveling-Wave Tubes using Transfer Matrices

### v1

@@ VIII.2 Two-Stage TWT
The two-stage TWT example was simulated using the PIC solver of CST Studio Suite with a simulation duration of 10 ns, approximately 29.9 million mesh cells, and approximately 2.27 million macroparticles. In this example, the input and output stages of the SWS each consist of $N_{\mathrm{c,1}}=N_{\mathrm{c,2}}=65$ unit cells. For this example, the peak gain of our generalized Pierce model (with perfect matching at the input and output ports) is also within approximately 0.9 dB of the peak gain found using PIC simulations, with the frequency of the peak gain for our model approximately 0.7 GHz higher than the peak gain frequency found from PIC simulations, as shown in Fig. 12a. The frequency of the peak gain calculated from LATTE was approximately 0.5 GHz below the peak gain obtained through PIC simulations, but the peak gain agrees within approximately 0.6 dB with our PIC simulations. Furthermore, the gain-versus-frequency profile of our model consistently overshoots the gain from PIC simulations at frequencies above 13 GHz, whereas LATTE consistently undershoots the gain from PIC simulations at the same frequencies. The state vector was again sampled at discrete positions along each TWT stage after being cascaded through transfer matrices and the sever like in the previous example. The power of the RF wave guided along the circuit is plotted vs distance at an operating frequency of 12 GHz, as shown in Fig. 12b. Plotting the RF power against the position-dependent attenuation profile of the SWS, we observe that the RF power increases up to about 70 mm from the input port, becomes extremely small at the sever gap, and then grows back with distance in the output stage of the TWT. Unlike the case of the single-stage TWT, the power of the state vector, $P(z)=\frac{1}{2}\mathrm{Re}\left(V(z)I^{*}(z)\right)$, for the two-stage TWT exhibits ripples along the length of the first stage with and without accounting for mismatches at the TWT ports; this is due to the reflections caused by the capacitive pi network of the sever shown in Fig. 10 that connects the two stages, as explained in Appendix D. As explained in the above subsection for the single-stage TWT, the apparent higher RF power at output port of the TWT when mismatches are present in Fig. 10 is due to our assertion that the accepted (i.e., input) RF power of the TWT is a constant $-10\,\mathrm{dBm}$, while the available RF power at the input of the TWT is larger when there is an input impedance mismatch. Furthermore, for the case of $-10\,\mathrm{dB}$ mismatches at the TWT ports, a gain ripple of approximately 5 dB near the center frequency is present. The gain ripple is more significant for the two-stage TWT example shown due to the higher gain and additional reflections at the sever that are not completely attenuated.
Figure 12: Same comparison as in Fig. 11, but for the two-stage TWT with a sever at 12 GHz.
@@ IX Comparison to LATTE
The TWT code LATTE requires the same frequency- and position-dependent inputs as our generalized Pierce model, which, aside from the electron beam parameters, are: interaction impedance, phase velocity, and the attenuation coefficient. However, LATTE differs from our model in that it is a Lagrangian-based model; it treats the electron beam as disks of charge, each interacting individually with the guided EM mode and other disks of charge (Wohlbier, Booske, and Dobson, 2002). Unlike our model, this feature allows LATTE to also simulate nonlinear effects such as intermodulation distortion, as well as large-signal TWT behavior. Many Eulerian- and Lagrangian-based TWT codes also exist which have been validated and studied in literature (Chong et al., 2005; Abe et al., 2000; Srivastava and Joshi, 1989; Chernin et al., 2001; Kory, 2001; Datta et al., 1998; Dialetis et al., 2009; Srivastava, 2003; Freund, 2000).
However, we believe that the way that LATTE models the sever gap in multi-stage TWTs is less accurate than our generalized Pierce model; LATTE ignores impedance mismatches at the sever gap and instead considers the sever gap as a continuous length of SWS. This approximation is valid only if the sever gap is negligibly small with respect to the guided wavelength and if there is sufficient attenuation to completely suppress reflections which occur at the sever gap. Furthermore, our transmission line-based model allows us to consider the effect of unavoidable mismatches at the input/output terminations of the TWT, in addition to mismatches that may occur at the sever gap. The code LATTE does not have such capabilities at this time. In the best case, any mismatches that are present on the TWT will only lead to gain ripple. In the worse case, the TWT may become an oscillator. The ability to calculate the effect of such mismatches in our generalized Pierce model makes it possible to obtain hot two-port parameters (i.e., the parameters relating equivalent circuit voltages and currents at the output of the tube to the input of the tube) for complex TWT designs and, in a future work, predict the small-signal stability like is done for two-port microwave transistor amplifiers in Ref. (Guillermo, 1984, Ch. 3) (in terms of the parameters $K$ and $\Delta$, there defined, which depend on active scattering parameters).
@@ X Conclusion
We have showcased a method for determining the small-signal gain of spatially inhomogeneous, lossy, and dispersive single-stage and multi-stage TWTs using a generalized Pierce model. The results of our model were compared against the Lagrangian TWT code LATTE and against full-wave PIC simulations in CST, with a peak gain that agrees within 1 dB and within 0.6 GHz with full-wave PIC simulations (when considering matched terminations at both ends of the TWT). Furthermore, we provide a Pierce-like model which uses the helix characteristic impedance, $Z_{\mathrm{c}}$, and a beam wave coupling factor, $a$, rather than the conventional interaction impedance, $Z_{\mathrm{P}}$. The imperfect but realistic helix sever in our multi-stage TWT model is represented using a capacitive pi-network. Furthermore, since our model is based on equivalent TLs and loads, we are able to consider how impedance mismatches at the input and output ports of the TWT lead to gain ripple.
@@ Appendix A Characteristic Impedance of Guided Modes in a Helix SWS
Following the work of (Kino and Paik, 1962; Paik, 1969), we determine the characteristic impedance of modes guided by a realistic helix SWS by approximating the tape helix as a lossless, anisotropically conducting tube, called the sheath helix, which was derived in (Sensiper, 1951). Using the sheath model, one may define the transmission line equations in terms of a voltage, $V$, defined as the potential difference between the surface of the sheath helix and the conducting wall which encloses the helix, and a current, $I$, which is defined as the longitudinal component of the current conducted by the sheath helix. The characteristic impedance of this sheath helix is also called the longitudinal characteristic impedance in (Pierce, 1950b; Paik, 1969; Henningsen, 1955). Paik also provides a transverse characteristic impedance definition in (Paik, 1969) which is proportional to the longitudinal impedance by the ratio of the group velocity to the phase velocity $v_{\mathrm{g}}/v_{\mathrm{c}}$, as explained in (Paik, 1969). In this paper, we wish to use the longitudinal characteristic impedance definition, since the it satisfies the power equation for forward voltage and current waves $P^{+}(z)=\frac{1}{2}Z_{\mathrm{c}}\left|I^{+}(z)\right|^{2}$, where the per-unit-length inductance and capacitance are calculated below as was done in (Kino and Paik, 1962; Paik, 1969). We will refer to the longitudinal characteristic impedance as “characteristic impedance” in the remainder of this appendix.
@@ Appendix B Alternative State Vector and System Matrix
As explained in the body of this paper and in Sec. A, the choice of voltage, current, and characteristic impedance definitions for a forward mode guided along the helix must satisfy the power relation $P^{+}(z)=\frac{1}{2}Z_{\mathrm{c}}\left|I^{+}(z)\right|^{2}$. Knowledge of this characteristic impedance is useful in the design of well-matched input or output ports on the TWT. However, in the original Pierce model, the characteristic impedance of the effective transmission line is the interaction impedance, $Z_{\mathrm{P}}$ (Pierce, 1947, 1950a, 1951). These two impedances are related in this paper by the coupling coefficient $a$ at each frequency. This coupling coefficient has also seen use in other works such as (Tamma and Capolino, 2014; Rouhi et al., 2021; Othman, Tamma, and Capolino, 2016; Othman et al., 2016; Abdelshafy et al., 2018; Figotin and Reyes, 2013; Figotin, 2021a; Abdelshafy et al., 2021; Figotin, 2021b; Abdelshafy, Capolino, and Figotin, 2022; Figotin, 2022, 2023).
As derived in the appendix of (Rouhi et al., 2024), the interaction impedance and characteristic impedance of the SWS are related through the frequency-dependent coupling coefficient $a$, as
@@
Using this relation between the helix characteristic impedance and interaction impedance, one can transform the transmission line voltage and current of the state vector in Eqn. (1) and system matrix in Eqn. (2) to be in terms of scaled TL quantities $V^{\prime}(z)=aV(z)$ and $I^{\prime}(z)=I(z)/a$ that maintain the average power definition $P=\frac{1}{2}\mathrm{Re}\left[V(z)I(z)^{*}\right]=\frac{1}{2}\mathrm{Re}\left[V^{\prime}(z)I^{\prime}(z)^{*}\right]$, where $a$ is the same coupling coefficient that relates interaction impedance to characteristic impedance. Looking at transformed forward propagating wave, the TL characteristc impedance is the same as the interaction (or Pierce) impedance, i.e., $V_{+}^{\prime}/I_{+}^{\prime}=Z_{\mathrm{P}}$. By making this transformation, the state vector may and system matrix are represented (with transmission line segment subscript $s$ suppressed) as
@@
We find that, with this transformation, the coupling coefficient is suppressed from the matrix and one can use $Z_{\mathrm{P}}$ as the characteristic impedance with the transformed TL parameters $V^{\prime}$ and $I^{\prime}$. This alternate formulation for the system matrix is useful, since the interaction impedance can be readily found for a realistic helix SWS using full-wave eigenmode simulations and postprocessing as in Eqn. (5), whereas it may be difficult to define and determine the equivalent voltage and current and the associated characteristic impedance in a realistic tape helix SWS. Furthermore, this alternate system matrix formulation more closely matches the original works by Pierce (Pierce, 1947, 1950a, 1951). The transformed system matrix $\mathbf{\underline{M}^{\prime}}$ may be used in place of the system matrix $\mathbf{\mathbf{\underline{M}}}$ shown in Eqn. (2) for segments of homogeneous transmission lines. Furthermore, the power gain and “hot” modal dispersion diagram (for homogeneous structures) of TWTs modeled using the transformed system matrix $\mathbf{\underline{M}^{\prime}}$ will be the same as those found using the original system matrix $\mathbf{\underline{M}}$, based on the fact that $P=\frac{1}{2}\mathrm{Re}\left[V(z)I(z)^{*}\right]=\frac{1}{2}\mathrm{Re}\left[V^{\prime}(z)I^{\prime}(z)^{*}\right]$. The only thing that differs in using $\boldsymbol{\Psi}(z)$ or $\boldsymbol{\Psi}^{\prime}(z)$ is that the matching terminations applied to the input and output of the SWS are either $Z_{\mathrm{c}}$ or $Z_{\mathrm{P}}$, respectively. This does not mean that the TWT ports should terminated in reality using one impedance or the other, it means that the definition of characteristic impedance of the equivalent transmisison line and source and load impedances must be consistent with the chosen voltage and current definitions. For helix-based SWSs, the source and load impedances $Z_{\mathrm{S}}$ and $Z_{\mathrm{L}}$ that matche with the SWS’ actual characteristic impedance are close to $Z_{\mathrm{c}}$ (using the sheath helix model to analytically approximate the characteristic impedance and modal dispersion of a tape helix, see Appendix A). When using the transformed TL parameters $V^{\prime}$ and $I^{\prime}$, the definition of impedance changes by a factor $a^{2}$ and, in this new basis, the transformed source and load impedances are $a^{2}Z_{\mathrm{S}}$ and $a^{2}Z_{\mathrm{L}}$, that match with the characteristic impedance $Z_{\mathrm{P}}$ of the SWS. The method described in this paper can be used also for unmatched loads, with the mismatch properly taken into account via the boundary conditions in Eqns. (15) or (16) for the single-stage or two-stage TWT, respectively, as shown in Sec. VIII.
@@
We note that in the body of this paper, we used the notation with $V$ and $I$ described in Sec. III, and the associated frequency-dependent characteristic impedance $Z_{\mathrm{c}}$ for the helix SWS. We also evaluated the frequency-dependent interaction impedance $Z_{\mathrm{P}}$ using results from the eigenmode solver of CST, as described in Sec. V and shown in Fig. 4. Using the calculated, frequency-dependent, characteristic impedance $Z_{\mathrm{c}}$ and the frequency-dependent interaction impedance $Z_{\mathrm{P}}$ from full-wave eigenmode simulations, we determined the frequency-dependent coupling coefficient $a$ through the relation in Eqn. 25.

### v2

@@ VIII.2 Two-Stage TWT
The two-stage TWT example was simulated using the PIC solver of CST Studio Suite with a simulation duration of 10 ns, approximately 29.9 million mesh cells, and approximately 2.27 million macroparticles. In this example, the input and output stages of the SWS each consist of $N_{\mathrm{c,1}}=N_{\mathrm{c,2}}=65$ unit cells. For this example, the peak gain of our generalized Pierce model (with perfect matching at the input and output ports) is also within approximately 0.9 dB of the peak gain found using PIC simulations, with the frequency of the peak gain for our model approximately 0.7 GHz higher than the peak gain frequency found from PIC simulations, as shown in Fig. 12a. The frequency of the peak gain calculated from LATTE was approximately 0.5 GHz below the peak gain obtained through PIC simulations, but the peak gain agrees within approximately 0.6 dB with our PIC simulations. Furthermore, the gain-versus-frequency profile of our model consistently overshoots the gain from PIC simulations at frequencies above 13 GHz, whereas LATTE consistently undershoots the gain from PIC simulations at the same frequencies. The state vector was again sampled at discrete positions along each TWT stage after being cascaded through transfer matrices and the sever like in the previous example. The power of the RF wave guided along the circuit is plotted vs distance at an operating frequency of 12 GHz, as shown in Fig. 12b. In addition, the approximate RF power is plotted using data from electric field probes in PIC simulations, using the method discussed in the previous subsection. Plotting the RF power against the position-dependent attenuation profile of the SWS, we observe that the RF power increases up to about 70 mm from the input port, becomes extremely small at the sever gap, and then grows back with distance in the output stage of the TWT. Unlike the case of the single-stage TWT, the power of the state vector, $P(z)=\frac{1}{2}\mathrm{Re}\left(V(z)I^{*}(z)\right)$, for the two-stage TWT exhibits ripples along the length of the first stage with and without accounting for mismatches at the TWT ports; this is due to the reflections caused by the capacitive pi network of the sever shown in Fig. 10 that connects the two stages, as explained in Appendix D. Gain ripples have also been observed in TWTs with multiple reflections in (Chernin et al., 2012). As explained in the above subsection for the single-stage TWT, the apparent higher RF power at output port of the TWT when mismatches are present in Fig. 10 is due to our assertion that the accepted (i.e., input) RF power of the TWT is a constant $-10\,\mathrm{dBm}$, while the available RF power at the input of the TWT is larger when there is an input impedance mismatch. Furthermore, for the case of $-10\,\mathrm{dB}$ mismatches at the TWT ports, a gain ripple of approximately 5 dB near the center frequency is present. The gain ripple is more significant for the two-stage TWT example shown due to the higher gain and additional reflections at the sever that are not completely attenuated. Gain ripple is generally undesireable for practical TWTs and can be mitigated in two ways: (a) by improving the match at the helix-to-coax or helix-to-waveguide transitions at the input and output port of the TWT with various geometries (Henningsen, 1955) or different dimensions, and (b) significantly increasing the attenuation leading into the sever gap region to further mitigate reflected waves.
Figure 12: (a) Gain vs frequency for a two-stage TWT with no mismatches on both ports ($\Gamma=0$, dashed blue line), and with a $-10\,\mathrm{dB}$ reflection on both ports ($\Gamma=0.1$, solid blue line) calculated using model described in this paper and compared to PIC and LATTE simulations. (b) Plot of attenuation coefficient and calculated RF power along the SWS versus longitudinal position at 12 GHz with and without a $-10\,\mathrm{dB}$ mismatch (dashed red and solid red lines, respectively) at the TWT ports using our generalized Pierce model, in addition to the approximate RF power calculated from field monitors in PIC simulations (black line with dots).
@@ IX Comparison to LATTE
The TWT code LATTE requires the same frequency- and position-dependent inputs as our generalized Pierce model, which, aside from the electron beam parameters, are: interaction impedance, phase velocity, and the attenuation coefficient. However, LATTE differs from our model in that it is a Lagrangian-based model; it treats the electron beam as disks of charge, each interacting individually with the guided EM mode and other disks of charge (Wohlbier, Booske, and Dobson, 2002). Unlike our model, this feature allows LATTE to also simulate nonlinear effects such as intermodulation distortion, as well as large-signal TWT behavior. Many Eulerian- and Lagrangian-based TWT codes also exist which have been validated and studied in literature (Chong et al., 2005; Abe et al., 2000; Srivastava and Joshi, 1989; Chernin et al., 2001; Kory, 2001; Datta et al., 1998; Dialetis et al., 2009; Srivastava, 2003; Freund, 2000). Furthermore, since the average voltage and current of the beam cannot vary in both the original and our generalized Pierce model, the beam cannot lose average power to the growing RF wave and net energy conservation is not possible in such a system. Due to this fact, our model is only useful in the small-signal regime. However, we also show a simple demonstration in Appendix H that, in the small-signal regime, the time-average power transfer per unit length between the beam and guided RF wave is nearly conserved under certain conditions.
However, in the small-signal regime, we believe that the way that LATTE models the sever gap in multi-stage TWTs is less accurate than our generalized Pierce model; LATTE ignores impedance mismatches at the sever gap and instead considers the sever gap as a continuous length of SWS. This approximation is valid only if the sever gap is negligibly small with respect to the guided wavelength and if there is sufficient attenuation to completely suppress reflections which occur at the sever gap. Furthermore, our transmission line-based model allows us to consider the effect of unavoidable mismatches at the input/output terminations of the TWT, in addition to mismatches that may occur at the sever gap. The code LATTE does not have such capabilities at this time. In the best case, any mismatches that are present on the TWT will only lead to gain ripple. In the worse case, the TWT may become an oscillator. The ability to calculate the effect of such mismatches in our generalized Pierce model makes it possible to obtain hot two-port parameters (i.e., the parameters relating equivalent circuit voltages and currents at the output of the tube to the input of the tube) for complex TWT designs and, in a future work, predict the small-signal stability like is done for two-port microwave transistor amplifiers in Ref. (Guillermo, 1984, Ch. 3) (in terms of the parameters $K$ and $\Delta$, there defined, which depend on active scattering parameters).
@@ X Conclusion
We have showcased a method for determining the small-signal gain of spatially inhomogeneous, lossy, and dispersive single-stage and multi-stage TWTs using a generalized Pierce model. The results of our model were compared against the Lagrangian TWT code LATTE and against full-wave PIC simulations in CST, with a peak gain that agrees within 1 dB and within 0.6 GHz with full-wave PIC simulations (when considering matched terminations at both ends of the TWT). Furthermore, our Pierce-like model uses the helix dispersive characteristic impedance, $Z_{\mathrm{c}}$, and a dispersive beam wave coupling factor, $a$, that depends on the interaction Pierce impedance, $Z_{\mathrm{P}}$. The imperfect but realistic helix sever in our multi-stage TWT model is represented using a capacitive pi-network. Furthermore, since our model is based on equivalent TLs and loads, we are able to consider how impedance mismatches at the input and output ports of the TWT lead to gain ripple.
@@ Appendix A Characteristic Impedance of Guided Modes in a Helix SWS
Following the work of (Kino and Paik, 1962; Paik, 1969), we determine the characteristic impedance of modes guided by a realistic helix SWS by approximating the tape helix as a lossless, anisotropically conducting tube, called the sheath helix, which was derived in (Sensiper, 1951). Using the sheath model, one may define the transmission line equations in terms of a voltage, $V$, defined as the potential difference between the surface of the sheath helix and the conducting wall which encloses the helix, and a current, $I$, which is defined as the longitudinal component of the current conducted by the sheath helix. The characteristic impedance of this sheath helix is also called the longitudinal characteristic impedance in (Paik, 1969; Henningsen, 1955). Paik also provides a transverse characteristic impedance definition in (Paik, 1969) which is proportional to the longitudinal impedance by the ratio of the group velocity to the phase velocity $v_{\mathrm{g}}/v_{\mathrm{c}}$, as explained in (Paik, 1969). In this paper, we wish to use the longitudinal characteristic impedance definition, since the it satisfies the power equation for forward voltage and current waves $P^{+}(z)=\frac{1}{2}Z_{\mathrm{c}}\left|I^{+}(z)\right|^{2}$, where the per-unit-length inductance and capacitance are calculated below as was done in (Kino and Paik, 1962; Paik, 1969). We will refer to the longitudinal characteristic impedance as “characteristic impedance” in the remainder of this appendix.
@@ Appendix B Alternative State Vector and System Matrix
As explained in the body of this paper and in Sec. A, the choice of voltage, current, and characteristic impedance definitions for a forward mode guided along the helix must satisfy the power relation $P=\frac{1}{2}\mathrm{Re}\left[V^{+}I^{+*}\right]$, where in the cold TL (i.e., without dependent current generator) one has $Z_{\mathrm{c}}=V^{+}/I^{+}$. Knowledge of this characteristic impedance is useful in the design of well-matched input or output ports on the TWT. However, in the original Pierce model, the characteristic impedance of the effective transmission line is the interaction impedance, $Z_{\mathrm{P}}$ (Pierce, 1947, 1951, 1950). These two impedances are related in this paper by the coupling coefficient $a$ at each frequency. This coupling coefficient has also seen use in other works such as (Tamma and Capolino, 2014; Rouhi et al., 2021; Othman, Tamma, and Capolino, 2016; Othman et al., 2016; Abdelshafy et al., 2018; Figotin and Reyes, 2013; Figotin, 2021a; Abdelshafy et al., 2021; Figotin, 2021b; Abdelshafy, Capolino, and Figotin, 2022; Figotin, 2022, 2023).
As derived in the appendix of (Rouhi et al., 2024), the interaction impedance and characteristic impedance of the cold SWS are related through the frequency-dependent coupling coefficient $a$, as
@@
Using this relation between the helix characteristic impedance and interaction impedance, one can transform the transmission line voltage and current of the state vector in Eqn. (1) and system matrix in Eqn. (2) to be in terms of scaled TL quantities $V^{\prime}(z)=aV(z)$ and $I^{\prime}(z)=I(z)/a$ that maintain the average power definition $P=\frac{1}{2}\mathrm{Re}\left[V(z)I(z)^{*}\right]=\frac{1}{2}\mathrm{Re}\left[V^{\prime}(z)I^{\prime}(z)^{*}\right]$, where $a$ is the same coupling coefficient that relates interaction impedance to characteristic impedance. Looking at transformed forward propagating wave, the TL characteristc impedance is the same as the interaction (or Pierce) impedance, i.e., $V^{\prime+}/I^{\prime+}=Z_{\mathrm{P}}$. By making this transformation, the state vector and system matrix are represented (with transmission line segment subscript $s$ suppressed) as
@@
We find that, with this transformation, the coupling coefficient is suppressed from the matrix and one can use $Z_{\mathrm{P}}$ as the characteristic impedance with the transformed TL parameters $V^{\prime}$ and $I^{\prime}$. This alternate formulation for the system matrix is useful, since the interaction impedance can be readily found for a realistic helix SWS using full-wave eigenmode simulations and postprocessing as in Eqn. (5), whereas it may be difficult to define and determine the equivalent voltage and current and the associated characteristic impedance in a realistic tape helix SWS. Furthermore, this alternate system matrix formulation more closely matches the original works by Pierce (Pierce, 1947, 1951, 1950). The transformed system matrix $\mathbf{\underline{M}^{\prime}}$ may be used in place of the system matrix $\mathbf{\mathbf{\underline{M}}}$ shown in Eqn. (2) for segments of homogeneous transmission lines. Furthermore, the power gain and “hot” modal dispersion diagram (for homogeneous structures) of TWTs modeled using the transformed system matrix $\mathbf{\underline{M}^{\prime}}$ will be the same as those found using the original system matrix $\mathbf{\underline{M}}$, based on the fact that $P=\frac{1}{2}\mathrm{Re}\left[V(z)I(z)^{*}\right]=\frac{1}{2}\mathrm{Re}\left[V^{\prime}(z)I^{\prime}(z)^{*}\right]$. The only thing that differs in using $\boldsymbol{\Psi}(z)$ or $\boldsymbol{\Psi}^{\prime}(z)$ is that the matching terminations applied to the input and output of the SWS are either $Z_{\mathrm{c}}$ or $Z_{\mathrm{P}}$, respectively. This does not mean that the TWT ports should terminated in reality using one impedance or the other, it means that the definition of characteristic impedance of the equivalent transmisison line and source and load impedances must be consistent with the chosen voltage and current definitions. For helix-based SWSs, the source and load impedances $Z_{\mathrm{S}}$ and $Z_{\mathrm{L}}$ that match the SWS’ actual characteristic impedance are close to $Z_{\mathrm{c}}$ (using the sheath helix model to analytically approximate the characteristic impedance and modal dispersion of a tape helix, see Appendix A). When using the transformed TL parameters $V^{\prime}$ and $I^{\prime}$, the definition of impedance changes by a factor $a^{2}$ and, in this new basis, the transformed source and load impedances are $a^{2}Z_{\mathrm{S}}$ and $a^{2}Z_{\mathrm{L}}$, that match with the characteristic impedance $Z_{\mathrm{P}}$ of the SWS. The method described in this paper can be used also for unmatched loads, with the mismatch properly taken into account via the boundary conditions in Eqns. (15) or (16) for the single-stage or two-stage TWT, respectively, as shown in Sec. VIII.
@@
We note that in the body of this paper, we used the notation with $V$ and $I$ described in Sec. III, and the associated frequency-dependent characteristic impedance $Z_{\mathrm{c}}$ for the helix SWS. We also evaluated the frequency-dependent interaction impedance $Z_{\mathrm{P}}$ using results from the eigenmode solver of CST Microwave Studio, as described in Sec. V and shown in Fig. 4. Using the calculated, frequency-dependent, characteristic impedance $Z_{\mathrm{c}}$ and the frequency-dependent interaction impedance $Z_{\mathrm{P}}$ from full-wave eigenmode simulations, we determined the frequency-dependent coupling coefficient $a$ through the relation in Eqn. 25.